OpenStudy (anonymous):

What is the value of n so that the expression x² + 12x + n is a perfect square trinomial? 6 36 72 144

Directrix (directrix):

Take half the coefficient of the linear x term and square it.

OpenStudy (anonymous):

ok but is it a, b, c,or d

Directrix (directrix):

12/2 = 6 6*6 = 36 x² + 12x + 36 = (x + 6)(x + 6) = (x+6)² which expands back to the perfect square trinomial.

Directrix (directrix):

The letter assigned to 36. n = 36

OpenStudy (anonymous):

thank you

Directrix (directrix):

Glad to help, @cupcakekiller1234